If a dress cost $23, and a retailer plans to sell it for $80, whats the dollar markup, percentage (of retail price) mark up, and percentage (of cost) mark up?

Dollar markup

Percentage (of retail price) markup

Percentage (of cost) markup
Answer #1

Dollar Mark up=Retail Price-Cost=80-23=$57

Percentage (of retail price) markup=Dollar Markup/Retail price=57/80=71.25%

Percentage (of cost) markup=Dollar Markup/Cost=57/23=247.83%